# Realistic overground gait transitions are not sharp but involve gradually changing walk-run mixtures as per energy optimality

### v1

@@ 4 Discussion
We have performed overground gait transition experiments over much longer distances than earlier and shown that humans use remarkably similar behavior to short distance bouts. Specifically, we find a gradually morphing walk-run transition regime, that involves walk-run mixtures, dominated by walking at lower speeds and by running at higher average speeds. These overground walk-run mixtures necessarily involve transitions from walking at a low speed to running at a high speed and vice versa, so perhaps the hysteretic behavior on a treadmill is a reflection of the overground strategy deployed in an unfamiliar treadmill setting.
@@
Aside from instantaneous energetics, gait transition from walking to running has been attributed to muscle force-velocity behavior [31], interlimb coordination variability [32], mechanical load or stress [33, 18], and cognitive or perceptual factors [34, 35]; see [36] for a review. However, none of these factors can show why there may be hystereses between the walk-to-run and the run-to-walk speeds on treadmills [37]. And importantly, we know it is unclear how any of these ‘instantaneous’ theories with the gradual walk-run-mixture-based gait transition regime behavior we have demonstrated in this manuscript. Predicting a walk-run mixture over a full bout as being optimal must necessarily require a theory that integrates some performance measure over the entire bout.
@@
We have considered one kind of overground gait transition, in which the task is traveling a given distance at a desired average sub-maximal speed. Another kind of ecological overground gait transition that might have been common in our evolutionary past is to be walking normally and then having to accelerate to a higher running speed to either chase prey or evade a predator. Piers et al [38] considered such a gait-transition task, but this is a substantially different task from that considered here from an energy optimality perspective, as it will be dominated by the cost of changing speeds and the cost of switching gaits [27, 28] – so we do not a priori expect identical gait transition speeds.
Future work could involve understanding what factors mediate the triggering of transitions between walking and running, as well as speed changes within a gait, based on the task requirements.
Acknowledgment
@@
This work was supported in part by NSF CMMI grant 1254842. Thanks to Mark Snartese and Max Donelan in suggesting the use of VBOX GPS over a less accurate GPS device, enabling the original study [1] as well as this current study.

### v2

@@ 4 Discussion
We have performed overground gait transition experiments over much longer distances than earlier [1] and shown that humans use remarkably similar behavior to short distance bouts. Specifically, we find a gradually morphing walk-to-run transition regime, that involves walk-run mixtures, dominated by walking at lower speeds and by running at higher average speeds. Further, in contrast to treadmill gait transition experiments, here, humans do not use walking up to some instantaneous speed and running above that speed. Instead, within each trial, the overground walk-run mixtures involve transitions from walking at a low speed directly to running at a high speed and vice versa, generally with a gap between these speeds. The avoidance of a range of intermediate speeds has been previously found in free-living animals such as ostriches, emus, horses and gnus [42, 8, 43, 44]. We might speculate that the hysteretic gait-transition behavior on a treadmill [18, 11, 12] could be an extrapolation of the overground strategy with such speed gaps, deployed to an unfamiliar treadmill setting.
@@
Aside from instantaneous energetics [5, 8], gait transition from walking to running has been attributed to muscle force-velocity behavior [45], interlimb coordination variability [46], mechanical load or stress [47, 18], cardiovascular loads [48], and cognitive or perceptual factors such as visual flow [49, 50], all examined on a treadmill; see [51] for a review. There is mixed evidence on whether cognitive load influences the walk to run gait transition [52, 53]. However, none of these instantaneous factors has been used to explain why there may be hystereses between the walk-to-run and the run-to-walk speeds on treadmills, while Diedrich and Warren [18, 54] provided a multiple-energy-minima-based explanation of such hysteresis. In contrast, gait transition mechanisms that are dynamical systems and stability-based [55, 18, 56, 57, 58] cannot immediately explain the natural overground transitions observed here as a function of average speed. Specifically, it is unclear how any of these ‘instantaneous’ theories can explain the gradual walk-run-mixture-based gait transition regime behavior we have demonstrated in this manuscript. Predicting a walk-run mixture over a full bout as being optimal must necessarily require a theory that integrates some performance measure over the entire bout, rather than making decisions instantaneously based on crossing a threshold — or at the very least, use a control policy that switches gait and speed based on the time and distance remaining. Such control strategies could be explored in future work. Some recent work has also pointed to usefulness of time-like costs or discounting to explain gait-speed choices, in addition to energy, but given such considerations do not apply here due to the precise time durations provided for each locomotor task [59, 37].
@@
We have considered one kind of overground gait transition, in which the task is traveling a given distance at a desired average sub-maximal speed. Another kind of ecological overground gait transition that might have been common in our evolutionary past is to be walking normally and then having to accelerate to a higher running speed to either chase prey or evade a predator. Piers et al [60] considered such an overground gait-transition task, but theirs is a substantially different task from that considered in this manuscript: from an energy optimality perspective, the task of Piers et al [60] will be dominated by the cost of changing speeds and the cost of switching gaits [21, 38] — so we do not a priori expect identical gait transition speeds.
The mathematical structure of the gait transition regime with walk-run mixtures is similar to that of phase transitions in materials, for instance, the existence of twinning microstructure (e.g., in an austenite-martensite boundary) or the co-existence of different phases of matter at a given temperature (e.g., water and vapor at room temperature, with the vapor fraction increasing with temperature). In a more complex mathematical model that was able to arbitrarily choose the locomotor patterns using trajectory optimization [27, 61, 2], a walking-like gait was selected at low speeds, a running-like gait was selected at high speeds, and the gait transition phase included a ‘flat’ portion of the energy landscape in which there were infinitely many gaits of equal energy cost in the simple approximation [61]. Our work has focused on a dynamics-free model of gait transition, but future work may involve performing long-timescale gait optimization with a biped model to obtain the optimal gait strategy at each moment [62]. Future work could also involve examining what factors mediate the triggering of transitions between walking and running, as well as speed changes within a gait, based on the task requirements. Finally, Long and Srinivasan [1] predicted the occurrence walk-run mixtures in small children (leg length 50-55 cm) walking next to adults, walk-trot mixtures in small dogs on a long leash walking next to a human, stand-walk or stand-run mixtures in small animals or even humans on sufficiently long treadmills, and in human 5-hour marathons, as all of these situations correspond to when such gait mixtures would be energy-optimal. Future studies could test each of these predictions via targeted experiments or observational studies.
Acknowledgment.
@@
This work was supported in part by NSF CMMI grant 1254842. We are grateful to Mark Snartese and Max Donelan in suggesting the use of VBOX GPS instead of a less accurate GPS device, enabling the original study [1] as well as this current study.
Competing interests.
We declare that we have no competing interests.
Data and code availability.
Data and code are available at this website for review purposes:
https://tinyurl.com/BakerWalkRunData and will be available publicly on Dryad upon publication.
Supplementary Figures
Figure S1: The histograms of GPS-derived speeds for each subject classified into walking (blue) and running (red) bouts. This data was pooled across subjects to produce the main manuscript figures 2c and 3a. Given the large distances involved, no subject performed all four average speed tasks, so some axes do not have histogram data.